9:50 am – 7:30 pm
- found and fixed bug in
- blead-back in of probe signal: probably a sign of not sufficiently sealed chamber. Try air drying sample and then sealing really tight.
- hybe2 stain initially very bright
- hybe2 not very strong at all in previously imaged loci — 405 damage to existing probes?
- non-imaged loci give much better stains
- it does seem the conventional images are still substantially dimmer than they started out. Maybe we should turn off the heater after the hybe
- rehybing probe 2 to confirm this difference.
- need to make more STORM buffer
- repeat BXC + F06
- start CC analysis of C09 (eve locus). Looks beautiful.